# Describe-and-Dissect: Interpreting Neurons in Vision Networks with Language Models

### v1

@@ Abstract
In this paper, we propose Describe-and-Dissect (DnD), a novel method to describe the roles of hidden neurons in vision networks. DnD utilizes recent advancements in multimodal deep learning to produce complex natural language descriptions, without the need for labeled training data or a predefined set of concepts to choose from. Additionally, DnD is training-free, meaning we don’t train any new models and can easily leverage more capable general purpose models in the future. We have conducted extensive qualitative and quantitative analysis to show that DnD outperforms prior work by providing higher quality neuron descriptions. Specifically, our method on average provides the highest quality labels and is more than 2$\times$ as likely to be selected as the best explanation for a neuron than the best baseline.
Machine Learning, ICML
@@ 1 Introduction
Previous methods have gained a deeper understanding of DNNs by examining the functionality (also known as concepts) of individual neurons [^1] We conform to prior works’ notation and use ”neuron” to describe a channel in CNNs.. This includes works based on manual inspection (Erhan et al., 2009; Zhou et al., 2014; Olah et al., 2020; Goh et al., 2021), which can provide high quality description at the cost of being very labor intensive. Alternatively, Network Dissection (Bau et al., 2017) automated this labeling process by creating the pixelwise labeled dataset, Broden, where fixed concept set labels serve as ground truth binary masks for corresponding image pixels. The dataset was then used to match neurons to a label from the concept set based on how similar their activation patterns and the concept maps were. While earlier works, such as Network Dissection, were restricted to an annotated dataset and a predetermined concept set, CLIP-Dissect (Oikarinen & Weng, 2023) offered a solution by no longer requiring labeled concept data, but still requires a predetermined concept set as input. By utilizing OpenAI’s CLIP model, CLIP-Dissect matches neurons to concepts based on their activations in response to images, allowing for a more flexible probing dataset and concept set compared to previous works.
@@
However, these methods still share a major limitation: Concepts detected by certain neurons, especially in intermediate layers, prove to be difficult to encapsulate using the simple, often single-word descriptions provided in a fixed concept set. MILAN (Hernandez et al., 2022) sought to enhance the quality of these neuron labels by providing generative descriptions, but their method requires training a new descriptions model from scratch to match human explanations on a dataset of neurons. This leads to their proposed method being more brittle and often performs poorly outside its training data.
@@
To overcome these limitations, we propose Describe-and-Dissect (abbreviated as DnD) in Section 3, a pipeline to dissect DNN by utilizing an image-to-text model to describe highly activating images for corresponding neurons. The descriptions are then semantically combined by a large language model, and finally refined with synthetic images to generate the final concept of a neuron. We conduct extensive qualitative and quantitative analysis in Sec 4 and show that Describe-and-Dissect outperforms prior work by providing high quality neuron descriptions. Specifically, we show that Describe-and-Dissect provides more complex and higher-quality descriptions (up to 2-4$\times$ better) of intermediate layer neurons than other contemporary methods in a large scale user study. Example descriptions from our method are displayed in Figure 1 and Appendix A.2. Additionally, we provide an interesting use-case study showing that these neuron descriptions can be used to find a good classifier for a class missing from the training set.
@@
Table 1: Comparison of existing automated neuron labeling methods and our Describe-and-Dissect (DnD). Green and boldfaced Yes or No indicates the desired property for a column. DnD has all the desired properties while existing work has some limitations.
@@ 2.1 Neuron Interpretability Methods
Network Dissection (Bau et al., 2017) is the first method developed to automatically describe individual neurons’ functionalities. The authors first defined the densely-annotated dataset Broden, denoted as $\mathcal{D}_{\textrm{Broden}}$, as a ground-truth concept mask. The dataset is composed of various images $x_{i}$, each labeled with concepts $c$ at the pixel-level. This forms a ground truth binary mask $L_{c}(x_{i})$ which is used to calculate the intersection over union (IoU) score between $L_{c}(x_{i})$ and the binary mask from the activations of the neuron $k$ over all images $x_{i}\in\mathcal{D}_{\textrm{Broden}}$, denoted $M_{k}(x_{i})$: $\textrm{IoU}_{k,c}=\frac{\sum_{x_{i}\in\mathcal{D}_{\textrm{Broden}}}M_{k}(x_{i})\cap L_{c}(x_{i})}{\sum_{x_{i}\in\mathcal{D}_{\textrm{Broden}}}M_{k}(x_{i})\cup L_{c}(x_{i})}.$ The concept $c$ is assigned to a neuron $k$ if $\textrm{IoU}_{k,c}>\eta$, where the threshold $\eta$ was set to 0.04. Intuitively, this method finds the labeled concept whose presence in the image is most closely correlated with the neuron having high activation. Extensions of Network Dissection were proposed by (Bau et al., 2020) and (Mu & Andreas, 2020).
@@
However, Network Dissection is limited by the need of concept annotation and the concept set is a closed set that may be hard to expand. To address these limitations, a recent work CLIP-Dissect (Oikarinen & Weng, 2023) utilizes OpenAI’s multimodal CLIP (Radford et al., 2021) model to describe neurons automatically without requiring annotated concept data. They leverage CLIP to score how similar each image in the probing dataset $\mathcal{D}_{probe}$ is to the concepts in a user-specified concept set to generate a concept activation matrix. To describe a neuron, they compare the activation pattern of said neuron to activations of different concepts on the probing data, and find the concept that is the closest match using a similarity function, such as softWPMI. Another very recent work FALCON (Kalibhat et al., 2023) uses a method similar to CLIP-Dissect but augments it via counterfactual images by finding inputs similar to highly activating images with low activation for the target neuron, and utilizing spatial information of activations via cropping. However, they solely rely on cropping the most salient regions within a probing image to filter spurious concepts that are loosely related to the ground truth functionality labels of neurons. This approach largely restrict their method to local concepts while overlooking holistic concepts within images, as also noted in (Kalibhat et al., 2023). Their approach is also limited to single word / set of words description that is unable to reach the complexity of natural language.
@@
MILAN (Hernandez et al., 2022) is a different approach to describe neurons using natural language descriptions in a generative fashion. Note that despite the concept sets in CLIP-Dissect and  FALCON being flexible and open, they cannot provide generative natural language descriptions like MILAN. The central idea of MILAN is to train an images-to-text model from scratch to describe the neuron’s role based on 15 most highly activating images. Specifically, it was trained on crowdsourced descriptions for 20,000 neurons from selected networks. MILAN can then generate natural language descriptions to new neurons by outputting descriptions that maximize the weighted pointwise mutual information (WPMI) between the description and the active image regions. One major limitation of MILAN is that the method require training a model to imitate human descriptions of image regions on relatively small training dataset, which may cause inconsistency and poor explanations further from training data. In contrast, our DnD is training-free, generative, and produces a higher quality of neuron descriptions as supported by our extensive experiments in Figure 1, Table 3, and Table 4. A detailed comparison between our method and the baseline methods is shown in Table 1.
@@ 2.2 Leveraging Large Pretrained models
The second model is GPT-3.5 Turbo, which is a transformer model developed by OpenAI for understanding and generating natural language. It provides increased performance from other contemporary models due to its vast training dataset and immense network size. We utilize GPT-3.5 Turbo for natural language processing and semantic summarization in the step 2 of our DnD. We use GPT-3.5 Turbo in this work as it’s one of the SOTAs in LLMs and cheap to use, but our method is compatible with other future and more advanced LLMs.
The third model is Stable Diffusion (Rombach et al., 2022), which is a text-to-image latent diffusion model (LDM) trained on a subset from the LAION-5B database (Schuhmann et al., 2022). By performing the diffusion process over the low dimensional latent space, Stable Diffusion is significantly more computationally efficient than other diffusion models. Due to its open availability, lower computational cost, and high performance, we employ Stable Diffusion for our image generation needs in the step 3 of DnD.
Figure 2: Overview of Describe-and-Dissect (DnD) algorithm. Given a Target model, it consists three important steps to identify the neuron concepts (e.g. ‘Swimming Shark‘ for neuron $n$).
@@ 3 Describe-and-Dissect: Methods
  • Step 1. Probing Set Augmentation: Augment the probing dataset with attention cropping to include both global and local concepts;
  • Step 2. Candidate Concept Generation: Generate initial concepts by describing highly activating images and subsequently summarize them into candidate concepts using GPT;
  • Step 3. Best Concept Selection: Generate new images based on candidate concepts and select the best concept based on neuron activations on these synthetic images with a scoring function.
An overview of Describe-and-Dissect (DnD) and these 3 steps are illustrated in Fig. 2.
@@ 3.1 Step 1: Probing Set Augmentations
Probing dataset $\mathcal{D}_{probe}$ is the set of of images we record neuron activations on before generating a description. As described in Section 2.1, one major limitation of (Kalibhat et al., 2023) is the restriction to local concepts while overlooking holistic concepts within images, while one limitation of (Oikarinen & Weng, 2023) is not incorporating the spatial activation information. Motivated by these limitations, DnD resolves these problems by augmenting the original probing dataset with a set of attention crops of the highest activating images from the original probing dataset. The attention crops can capture the spatial information of the activations and we name this set as $\mathcal{D}_{cropped}$, shown in Fig. 2. We discuss the implementation details of our attention cropping procedure in Appendix A.1.1 and an perform an ablation study of its effects in Appendix A.4.1.
@@ 3.2 Step 2: Candidate Concept Generation
The output of Step 2B is a set of $N$ descriptions which we call ”candidate concepts”. We denote this set as $T=\{T_{1},...,T_{N}\}$. For the purposes of our experiments, we generate $N=5$ candidate concepts unless otherwise mentioned. The exact prompt used for GPT summarization is shown in Appendix A.1.2.
@@ 3.3 Step 3: Best Concept Selection
The last crucial component of DnD is concept selection, which selects the concept from the set of candidate concepts $T$ that is most correlated to the activating images of a neuron. We first use the Stable Diffusion model (Rombach et al., 2022) from Hugging Face to generate images for each concept $T_{j\in[N]}$. Generating new images is important as it allows us to differentiate between neurons truly detecting a concept or just spurious correlations in the probing data. The resulting set of images is then fed through the target model again to record the activations of a target neuron on the new images. Finally, the candidate concepts are ranked using a concept scoring model, as discussed below in section 3.4.
@@
[…]
In simpler terms, the intuition behind this algorithm is that if a neuron $n$ encodes for a concept $c$, then the images generated to encapsulate that concept $c$ should cause the neuron $n$ to activate highly. While we only experiment with Best Concept selection within the DnD framework, it can be independently applied with other methods like (Bau et al., 2017; Hernandez et al., 2022; Oikarinen & Weng, 2023) to select the best concept out of their top-k best descriptions, which is another benefit of our proposed method.
@@ 3.4 Scoring Function
An integral part of Best Concept Selection is the scoring function. This function takes in the information provided by the activations of the candidate concepts’ synthetic images and computes a score to select the best candidate concept. In this paper we propose 3 different scoring functions and a combination of two of them: Mean, TopK Squared, Image Products, and TopK Squared + Image Products. See Sections A.1.3 for the details behind each of these functions. In section A.1.4, we compare the different functions and observe that all perform similarly well, showcasing our algorithm works robustly with different options. We use the TopK Squared + Image Products scoring function for all experiments unless otherwise specified.
@@ 4 Experiments
In this section, we present extensive qualitative and quantitative analysis to show that DnD outperforms prior works by providing higher quality neuron descriptions. For fair comparison, we follow the setup in prior works to run our algorithm on the following two networks: ResNet-50 and ResNet-18 (He et al., 2016) trained on ImageNet (Russakovsky et al., 2015) and Place365 (Zhou et al., 2016) respectively. In section 4.1, we qualitatively analyze DnD along with other methods on random neurons and show that our method provides good descriptions on these examples. Next in section 4.2 we quantitatively show that DnD yields superior results to comparable methods. In section 4.3, we show that our method outperforms existing neuron description methods in large scale crowdsourced studies. In section 4.4, we demonstrate a use case of DnD and provide another means of quantitative comparison with similar methods by utilizing Describe-and-Dissect to find classifiers for new concepts. Finally in section 4.5 we study the importance of each step in our pipeline through extensive ablation studies.

### v2

# Interpreting Neurons in Deep Vision Networks with Language Models

@@ Abstract
In this paper, we propose Describe-and-Dissect (DnD), a novel method to describe the roles of hidden neurons in vision networks. DnD utilizes recent advancements in multimodal deep learning to produce complex natural language descriptions, without the need for labeled training data or a predefined set of concepts to choose from. Additionally, DnD is training-free, meaning we don’t train any new models and can easily leverage more capable general purpose models in the future. We have conducted extensive qualitative and quantitative analysis to show that DnD outperforms prior work by providing higher quality neuron descriptions. Specifically, our method on average provides the highest quality labels and is more than 2$\times$ as likely to be selected as the best explanation for a neuron than the best baseline. Finally, we present a use case providing critical insights into land cover prediction models for sustainability applications. Our code and data are available at https://github.com/Trustworthy-ML-Lab/Describe-and-Dissect.
**footnotetext: N. Bai and R. Iyer contributed equally to this work. Major work done in Summer outreach program at UC San Diego.
@@ 1 Introduction
Previous methods have gained a deeper understanding of DNNs by examining the functionality (also known as concepts) of individual neurons [^1] We conform to prior works’ notation and use ”neuron” to describe a channel in CNNs.. This includes works based on manual inspection (Erhan et al., 2009; Zhou et al., 2015; Olah et al., 2020; Goh et al., 2021), which can provide high quality description at the cost of being very labor intensive. Alternatively, Network Dissection (Bau et al., 2017) automated this labeling process by creating the pixelwise labeled dataset, Broden, where fixed concept set labels serve as ground truth binary masks for corresponding image pixels. The dataset was then used to match neurons to a label from the concept set based on how similar their activation patterns and the concept maps were. While earlier works, such as Network Dissection, were restricted to an annotated dataset and a predetermined concept set, CLIP-Dissect (Oikarinen & Weng, 2023) offered a solution by no longer requiring labeled concept data, but still requires a predetermined concept set as input. By utilizing OpenAI’s CLIP model, CLIP-Dissect matches neurons to concepts based on their activations in response to images, allowing for a more flexible probing dataset and concept set compared to previous works.
@@
However, these methods still share a major limitation: Concepts detected by certain neurons, especially in intermediate layers, prove to be difficult to encapsulate using the simple, often single-word descriptions provided in a fixed concept set. MILAN (Hernandez et al., 2022) sought to enhance the quality of these neuron labels by providing generative descriptions, but their method requires training a new description model from scratch to match human explanations on a dataset of neurons. This leads to their proposed method being more brittle and often performs poorly outside its training data.
@@
To overcome these limitations, we propose Describe-and-Dissect (abbreviated as DnD) in Section 3, a pipeline to dissect DNNs by utilizing an image-to-text model to describe highly activating images for corresponding neurons. The descriptions are then semantically combined by a large language model, and finally refined with synthetic images to generate the final concept of a neuron. We conduct extensive qualitative and quantitative analysis in Section 4 and show that Describe-and-Dissect outperforms prior work by providing high quality neuron descriptions. Specifically, we show that Describe-and-Dissect provides more complex and higher-quality descriptions (up to 2-4$\times$ better) on intermediate layer neurons than other contemporary methods in a large scale user study. Example descriptions from our method are displayed in Figure 1. Additionally, we present a use-case study demonstrating DnD’s ability to interpret and improve upon current sustainability models in Section 5.
@@
Table 1: Comparison of existing automated neuron labeling methods and our method Describe-and-Dissect (DnD). Green and boldfaced Yes or No indicates the desired property for a column. DnD has all the desired properties while existing work has some limitations.
@@ 2.1 Neuron Interpretability Methods
Network Dissection (Bau et al., 2017) is the first method developed to automatically describe individual neurons’ functionalities. The authors first defined the densely-annotated dataset Broden, denoted as $\mathcal{D}_{\textrm{Broden}}$, as a ground-truth concept mask. The dataset is composed of various images $x_{i}$, each labeled with concepts $c$ at the pixel-level. This forms a ground truth binary mask $L_{c}(x_{i})$ which is used to calculate the intersection over union (IoU) score between $L_{c}(x_{i})$ and the binary mask from the activations of the neuron $k$ over all images $x_{i}\in\mathcal{D}_{\textrm{Broden}}$, denoted $M_{k}(x_{i})$: $\textrm{IoU}_{k,c}=\frac{\sum_{x_{i}\in\mathcal{D}_{\textrm{Broden}}}M_{k}(x_{i})\cap L_{c}(x_{i})}{\sum_{x_{i}\in\mathcal{D}_{\textrm{Broden}}}M_{k}(x_{i})\cup L_{c}(x_{i})}.$ The concept $c$ is assigned to a neuron $k$ if $\textrm{IoU}_{k,c}>\eta$, where the threshold $\eta$ was set to 0.04. Intuitively, this method finds the labeled concept whose presence in the image is most closely correlated with the neuron having high activation. Extensions of Network Dissection were proposed by Bau et al. (2020) and Mu & Andreas (2020).
@@
However, Network Dissection is limited by the need of concept annotation and the concept set is a closed set that may be hard to expand. To address these limitations, a recent work CLIP-Dissect (Oikarinen & Weng, 2023) utilizes OpenAI’s multimodal CLIP (Radford et al., 2021) model to describe neurons automatically without requiring annotated concept data. They leverage CLIP to score how similar each image in the probing dataset $\mathcal{D}_{probe}$ is to the concepts in a user-specified concept set to generate a concept activation matrix. To describe a neuron, they compare the activation pattern of said neuron to activations of different concepts on the probing data, and find the concept that is the closest match using a similarity function, such as softWPMI. Another very recent work FALCON (Kalibhat et al., 2023) uses a method similar to CLIP-Dissect but augments it via counterfactual images by finding inputs similar to highly activating images with low activation for the target neuron, and utilizing spatial information of activations via cropping. However, they solely rely on cropping the most salient regions within a probing image to filter spurious concepts that are loosely related to the ground truth functionality labels of neurons. This approach largely restrict their method to local concepts while overlooking holistic concepts within images, as noted by Kalibhat et al. (2023). Their approach is also limited to single word/set of words descriptions that are unable to reach the complexity of natural language.
@@
On the other hand, MILAN (Hernandez et al., 2022) is a different approach to describe neurons using natural language descriptions in a generative fashion. Note that despite the concept sets in CLIP-Dissect and FALCON being flexible and open, they cannot provide generative natural language descriptions like MILAN. The central idea of MILAN is to train an image-to-text model from scratch to describe the neuron’s role based on 15 most highly activating images. Specifically, it was trained on crowdsourced descriptions for 20,000 neurons from selected networks. MILAN can then generate natural language descriptions to new neurons by outputting descriptions that maximize the weighted pointwise mutual information (WPMI) between the description and the active image regions. One major limitation of MILAN is that the method require training a model to imitate human descriptions of image regions on relatively small training dataset, which may cause inconsistency and poor explanations further from training data. In contrast, our DnD is training-free, generative, and produces a higher quality of neuron descriptions as supported by our extensive experiments in Figure 1 and Table 3. A detailed comparison between our method and the baseline methods is shown in Table LABEL:tab:method_strengths.
@@ 2.2 Leveraging Large Pretrained models
The second model is GPT-3.5 Turbo, which is a transformer model developed by OpenAI for understanding and generating natural language. It provides increased performance from other contemporary models due to its vast training dataset and immense network size. We utilize GPT-3.5 Turbo for natural language processing and semantic summarization in the Step 2 of our DnD. We use GPT-3.5 Turbo in this work as it’s one of the SOTAs in LLMs and cheap to use, but our method is compatible with other future and more advanced LLMs. We provide a quantitative comparison analyzing the effect of GPT-3.5 Turbo, GPT-4.0, and LLaMA2 on DnD’s label quality in Appendix A.4.5 as well as evaluation on cost and usage limitations for each model.
The third model is Stable Diffusion (Rombach et al., 2022), which is a text-to-image latent diffusion model (LDM) trained on a subset from the LAION-5B database (Schuhmann et al., 2022). By performing the diffusion process over the low dimensional latent space, Stable Diffusion is significantly more computationally efficient than other diffusion models, such as DALLE (Ramesh et al., 2021). Due to its open availability, lower computational cost, and high performance, we employ Stable Diffusion for our image generation needs in the Step 3 of DnD.
@@ 3 Describe-and-Dissect: Methods
  • Step 1. Probing Set Augmentation: Augment the probing dataset with attention cropping to include both global and local concepts. This helps better describe localized neuron activations.
  • Step 2. Candidate Concept Generation: Generate initial concepts by describing highly activating images and subsequently summarize them into candidate concepts using GPT. This is the main step that generates a set of possible explanations.
  • Step 3. Best Concept Selection: Generate new images based on candidate concepts and select the best concept based on neuron activations on these synthetic images with a scoring function. This refines the description by selecting the most accurate neuron description from descriptions produced in Step 2.
An overview of Describe-and-Dissect (DnD) and these 3 steps are illustrated in Figure  2.
Figure 2: Overview of Describe-and-Dissect (DnD) algorithm. For a given target model, DnD consists of three important steps to identify neuron concepts (e.g. ‘Swimming Shark‘ for neuron $n$).
@@ 3.1 Step 1: Probing Set Augmentations
Probing dataset $\mathcal{D}_{probe}$ is the set of images we record neuron activations on before generating a description. As described in Section 2.1, one major limitation of Kalibhat et al. (2023) is the restriction to local concepts while overlooking holistic concepts within images, while one limitation of Oikarinen & Weng (2023) is not incorporating the spatial activation information. Motivated by these limitations, DnD resolves these problems by augmenting the original probing dataset with a set of attention crops of the highest activating images from the original probing dataset. The attention crops can capture the spatial information of the activations and we name this set as $\mathcal{D}_{cropped}$, shown in Figure  2. We discuss the implementation details of our attention cropping procedure in Appendix A.1.1 and perform an ablation study on its effects in Appendix A.4.1.
@@ 3.2 Step 2: Candidate Concept Generation
The output of Step 2B is a set of $N$ descriptions which we call "candidate concepts". We denote this set as $T=\{T_{1},...,T_{N}\}$. For the purposes of our experiments, we generate $N=5$ candidate concepts unless otherwise mentioned. The exact prompt used for GPT summarization is shown in Appendix A.1.3.
@@ 3.3 Step 3: Best Concept Selection
The last crucial component of DnD is concept selection, which selects the concept from the set of candidate concepts $T$ that is most correlated to the activating images of a neuron. We first use the Stable Diffusion model (Rombach et al., 2022) from Hugging Face to generate images for each concept $T_{j\in[N]}$. Generating new images is important as it allows us to differentiate between neurons truly detecting a concept or just spurious correlations in the probing data. The resulting set of images is then fed through the target model again to record the activations of a target neuron on the new images. Finally, the candidate concepts are ranked using a concept scoring function, as discussed in Section 3.4.
@@
    Assign scores to each concept. The scoring function $\textit{score}(\mathcal{H}_{j})$ assigns a score to a concept using the rankings of the concept’s generated images, and potential additional information. The concept with the best (highest) score in $T$ is selected as the concept label for the neuron. Concept scoring functions are discussed below in Section 3.4.
In simpler terms, the intuition behind this algorithm is that if a neuron $n$ encodes for a concept $c$, then the images generated to encapsulate that concept $c$ should cause the neuron $n$ to activate highly. While we only experiment with Best Concept selection within the DnD framework, it can be independently applied with other methods like Bau et al. (2017); Hernandez et al. (2022); Oikarinen & Weng (2023) to select the best concept out of their top-k best descriptions, which is another benefit of our proposed method. DnD can also be used without Best Concept Selection (Step 3) to reduce computational costs.
@@ 3.4 Scoring Function
For a given neuron, we use a scoring function to rate candidate concept accuracy during Best Concept Selection (Step 3). Simple metrics such as mean are heavily prone to outliers that result in skewed predictions so we propose a scoring function that weights the average rank of top activating images mapping to a candidate concept.
$score(R_{j},I,\mathcal{D}_{j}^{t})=(N-\text{Rank}(R_{j}))\cdot E(I,\mathcal{D}_{j}^{t})$
Here, the average rank of images for candidate concept $j$, $\forall j\in\{1,...,N\}$, is denoted $R_{j}$ and $\text{Rank}(R_{j})$ sorts the average rank $R_{j}$ of each candidate concept in increasing order. $E(I,\mathcal{D}_{j}^{t})$ computes the average cosine similarity between image embeddings of $\mathcal{D}_{j}^{t}$ and $I$ using CLIP-ViT-B/16 (Radford et al., 2021), with $\mathcal{D}_{j}^{t}\subset\mathcal{D}_{j}$ for $t$ highest activating images (for our experiments, we use $t=10$). In practice, $R_{j}$ is computed as the square of the ranks in top $\beta=5$ ranking images for better differentiation between scores, $R_{j}=\{(R_{j}^{i})^{2};i\leq\beta\}$. Sections A.1.4 the details specifics behind the function. In Section A.1.5, we compare between various functions and show our algorithm works robustly with different options.
@@ 4 Experiments
In this section, we present extensive qualitative and quantitative analysis to show that DnD outperforms prior works by providing higher quality neuron descriptions. For fair comparison, we follow the setup in prior works to run our algorithm on the following two networks: ResNet-50 and ResNet-18 (He et al., 2016) trained on ImageNet (Russakovsky et al., 2015) and Place365 (Zhou et al., 2016) respectively. In Section 4.1, we qualitatively analyze DnD along with other methods on random neurons and show that our method provides good descriptions on these examples. Next in Section 4.2 we quantitatively show that DnD yields superior results to comparable methods. In Section 4.3, we show that our method outperforms existing neuron description methods in large scale crowdsourced studies. Finally in Section 4.4 we study the importance of critical steps in our pipeline by ablating away Generative Image Captioning (Step 2A) and Concept Selection (Step 3). Supplementary results are presented in the appendix, including method details in Section A.1, additional qualitative examples in Section A.2, comparison to MILANNOTATIONS in Section A.3, extensive ablation studies on each step of the DnD framework in Section A.4, an additional use case of DnD as an OOD classifier in Section A.6, and the capability to describe polysemantic neurons by producing multiple labels in Section A.7.